i raised to i?

i raised to i?

Grade:10

2 Answers

swapan jain
12 years ago

i=e^i(pie)/2,

i^i=e^((i*i)pie/2)

    =e^(-pie/2)
there can be many more answers.increase the angle by 2 pie. 

Chetan Mandayam Nayakar
12 years ago

let exp(.)≡ e(.),ii =exp(i((π/2)+2nπ))i =e-((π/2)+2nπ), where 'n' is ANY INTEGER